A depth of 10m of water exerts the same amount of pressure as the entire Earth's atmosphere which is~120km thick.Suggest why

1
Expert's answer
2022-02-07T08:43:05-0500

Given:

"h_{\rm water}=10\:\rm m"

"\rho_{\rm water}=1000\:\rm kg/m^3"

"h_{\rm air}=120\:000\:\rm m"

"\rho_{\rm air}=0.0833\:\rm kg/m^3\; (mean\; density)"


The hydrostatic pressure is given by

"p=\rho g h"

We get

"p_{\rm water}=1000*9.8*10=98\: 000\:\rm Pa"

"p_{\rm air}=0.0833*9.8*120\:000=98\: 000\:\rm Pa"